# From Insights to Actions: The Impact of Interpretability and Analysis Research on NLP

### v1

@@ 7 Main takeaways and discussion
In §4, we saw that IA research plays a central role in NLP and researchers build on findings from IA work in their research, regardless of whether they work on IA themselves or not. In section §5, we saw that NLP researchers and practitioners perceive IA work to be important for progress in NLP, and multiple subfields. They also find it important for their own work for a variety of reasons, regardless of whether they work on IA themselves. Finally, we took a closer look at the most influential IA papers in §6 and found that many novel methods are proposed based on IA findings and highly influenced by them, for various areas, in particular, work on reasoning, factual knowledge, and bias. All these findings present a very positive view of IA research and its role within NLP in the past and the present. In the remainder of this section, we turn to the future of IA research.
@@
Based on our findings, we make the following recommendations:
Going forward, IA researchers should: 1. Think more about the big picture 2. Strive for more actionable work 3. Center humans in your work 4. Work towards standardized, robust methods
Big-picture thinking involves working towards general truths about model architectures or behaviors, rather than model-specific results. Actionable work requires thinking about how an IA finding can propel new ways of building/using NLP systems, rather than being merely descriptive. Centering humans entails evaluation with realistic and relevant data and tasks, and performing user studies and human evaluation. Human-centered IA work can also be enhanced through interdisciplinary reading and collaboration. Finally, we urgently need to build consensus on using and evaluating IA methods. Rigorous, well-motivated methods (e.g., using causality) are critical, rather than correlative evidence that may not be correct or faithful.
@@ 8 Conclusion
The starting point of our analysis are all papers published at ACL and EMNLP. Although these are the most cited *CL venues (Mohammad, 2020), our analysis excludes several other big NLP venues, including EACL, NAACL, AACL, TACL, and workshops, including BlackboxNLP, which focuses on IA work. Additionally, given the growing interest in NLP, and in particular, LLMs, from the broader machine learning community, there is an increasing number of IA papers published at machine learning conferences such as ICLR, NeurIPS, and ICML, which we also do not consider in our analyses. Similarly, a vast amount of work on mechanistic interpretability has been published as articles (e.g., on LessWrong [^6] https://www.lesswrong.com/ and the AI Alignment Forum [^7] https://www.alignmentforum.org/), and blog posts (e.g., by Anthropic [^8] https://www.anthropic.com/). Therefore, there is a risk that our analysis misses potentially influential IA work published at these venues.
@@
Our analysis focuses on papers published between 2018 and 2024. Our results thus represent a snapshot in time on the scale of research in NLP, where models and methods come and go. The time period that we look at is dominated by transformer-based language models, and a paradigm of using large, general-purpose pre-trained models for many tasks, and thus many IA papers focus on studying these. Understanding this as the context of our analysis and results is important, as they may look completely different in a time period where the most popular models are different or the most popular IA methods are different. This also means that our results cannot speak to the impact of today’s IA work as its true impact might only become clear in the future.
@@
Although our use of citations is an important component of how we quantify impact in this paper, we do not consider citational context or distinguish between types of citations. However, papers can cited for a number of reasons (Bornmann and Daniel, 2008), not all positive and not all having to do with the conventions of scholarly publishing (Bornmann and Daniel, 2008; Zhu et al., 2015; Bornmann and Marx, 2012).
Limitations of our survey
Although we took steps to get a large number and diversity of survey responses, and we ensured a minimum of 10 respondents per bucket when reporting disaggregated results, the 138 responses we received may not be representative of the field as a whole. In particular, full professors (N=5, at various career stages), and industry practitioners who are not researchers (N=1) were somewhat underrepresented in our responses, indicating that our results focus more on research impact rather than impact on industry applications, and are overwhelmingly shaped by PhD students (41.3% of respondents), whose interests, incentives, and assessment of impact are sure to be different from respondents at other career stages.
Some respondents brought up the following concerns: one respondent felt our definition of IA was too broad for their taste, but our inclusion of interpretability and analysis was by design (see Section 3). Another respondent noted that we defined IA but not what we meant by “progress,” which was also by design, as we did not want to impose a normative definition of progress on our respondents but rather, get at their own intuitions, regardless of how they might define progress. Finally, one respondent complained that our questions about the usefulness of IA (to various subfields, on one’s own research, etc.) were framed in absolute rather than relative terms, and that just because IA research has some positive impact on our understanding doesn’t mean that it is the best option to pursue given limited time and resources. This paper presents views of absolute and relative impact via the survey and citation graph analyses, for a holistic view of IA research that also allows for it to have value for its own sake. Ultimately, we believe that a view of “optimal” impact compared to other options lies in the eye of the beholder, and is one (but not the only) way of interpreting our results.
@@ Acknowledgments
Appendix A Related work
The increasing number of IA publications during the last few years has resulted in several survey or position papers that critically discuss existing work, identify common patterns, and provide suggestions for how to go forward. Lipton (2018) critically question common motivations behind interpretability and the lack of definitions in the field. We follow their recommendation and provide a definition of what we consider interpretability and analysis research in §2. Belinkov and Glass (2019) summarize trends in early IA work and discuss recommendations for how to overcome limitations of IA research. Similar to our work, they recommend that future work should think about better ways to evaluate IA research and findings. Rogers et al. (2020) survey and synthesize IA work on BERTology, a subfield of IA work that focuses on encoder-only language models. Rauker et al. (2023) survey a large number of papers that study the internals of language models (transparency), and discuss key challenges in the field. Similar to our work, they also argue for better ways of evaluating IA methods, as well as more actionability and grounding in real-world applications. More recently, Madsen et al. (2024) discuss two prominent trends in interpretability research (post-hoc explanations and intrinsic interpretability) and argue that interpretability (“the study of explaining models in understandable terms to humans”) needs a new paradigm, centered around faithfulness.
Several other works study citational patterns and trends within the broader NLP community. Mohammad (2020) uses citations to measure the impact of NLP publications indexed by the ACL Anthology. Similar to our approach, they compare how well papers from different areas within NLP are cited, and use citation statistics to draw conclusions about the impact of different subfields within NLP. Singh et al. (2023) consider citations as an indicator for how widely the community is reading. They study temporal citations trends and reveal that a majority of cited papers fall within a five year time period before publication of the citing work, demonstrating a recency bias in citation behavior. Jacovi (2023) uses Semantic Scholar to curate a large number of papers focusing on explainability, studying citation trends in the field based on this collection. Wahle et al. (2023) analyze the influence between NLP and other fields over the years. Also using Semantic Scholar, they rely on citations to conclude that NLP has become more insular over time.
Another set of related papers surveys the NLP community for their perceptions and opinions, a method we also use. Gururaja et al. (2023), for example, focus on paradigm shifts and study factors that shape NLP as a field. They conduct interviews with NLP researchers and experts and gather their opinions on critical trends and patterns that emerge in the field. Pramanick et al. (2023) also focus on paradigm shifts and impact, but from a diachronic perspective. They provide a novel framework to study the evolution of research topics within a field to establish what drives research in NLP across time. They find that tasks and methods have a bigger impact on the field than metrics do.
Lastly, there are several related works in the scientometrics literature that study and compare the impact of research using the same metrics as we do: Chacon et al. (2020) apply the citation success index to compare sub-fields in physics, and Leydesdorff (2007) propose the use of Betweenness Centrality as a measure of the interdisciplinarity of journals.
Appendix B Citation graph details
@@
Phonology, Morphology and 72
Word Segmentation
@@
Unsupervised and Weakly- 17
Supervised Methods in NLP
@@ B.1 Sanity checks
Figure 8: Betweenness centralities versus citation counts for papers in ACL and EMNLP since 2020.
@@
Appendix C Survey details
@@
C.1 Ethical considerations
Our survey involved research with human participants, thus we report the full text of the survey below, and information about recruitment in Section 3. We determined there to be a negligible risk of harms from participating in our survey, as it contains no offensive or harmful content. As shown in the full survey below, we describe our study objectives and remind respondents that filling out the survey is completely voluntary. We then explicitly ask for their consent to participate, and obtain consent from all 138 survey respondents. For respondents who may not have completed the survey, no data was collected. In lieu of financial compensation, we offered survey respondents the optional opportunity to provide their name or an alias that we would mention in the acknowledgements of any future paper we write with the survey results. To protect respondent privacy and confidentiality, we will not release the original survey responses in full, but only release high-level statistics, annotations from our qualitative coding, and select non-identifying examples in Section 7.
C.2 Full survey
@@ 9. Do you think model analysis and interpretability research is important, and if so, why?
Appendix D Qualitative coding
@@
Appendix E Additional results
@@
Figure 10: Betweenness centrality of ACL and EMNLP papers since 2020 by track. Lines at the middle of the box represent the medians, but some tracks have their median at 0.
@@
Figure 11: Percentage of references to IA papers according to our classifiers prediction.
Figure 12: Ratio of intra-track citations according to the predictions of our classifier.

### v2

@@ 7 Main takeaways and discussion
In §4, we saw that IA research plays a central role in NLP and researchers build on findings from IA work in their research, regardless of whether or not they work on IA themselves. In §5, we saw that NLP researchers and practitioners perceive IA work to be important for progress in NLP, and multiple subfields. They also find it important for their own work for a variety of reasons, regardless of whether they work on IA themselves. Finally, we took a closer look at the most influential IA papers in §6 and found that many novel methods are proposed based on IA findings and highly influenced by them, for various areas—in particular, work on reasoning, factual knowledge, and bias. All these findings present a very positive view of IA research and its role within NLP in the past and the present. In the remainder of this section, we turn to the future of IA research.
@@
Based on our findings, we make the following recommendations for IA work:
Going forward, IA researchers should: 1. Think more about the big picture 2. Strive for more actionable work 3. Center humans in their research 4. Work towards standardized, robust methods
Concretely, big-picture thinking (1) involves working towards general truths about model architectures or behaviors, rather than model-specific results. Future work should try to synthesize existing strands of research to unify their findings and viewpoints. An example of what this might look like outside IA research is He et al. (2022).
Actionable work (2) requires thinking about how an IA finding can propel new ways of building/using NLP systems, rather than merely being descriptive. More specific examples of this include research that uses interpretability findings to, e.g., improve the fairness of NLP systems, or make NLP models more efficient and robust.
Centering humans (3) entails evaluation with realistic, relevant data and tasks, and performing user studies and human evaluation. Human-centered IA work can also be enhanced through interdisciplinary reading and collaboration. An example for research that falls under this category is Ivanova et al. (2024), which proposes a cognition-inspired framework for evaluating LLM world knowledge.
Finally, we urgently need to build consensus on using and evaluating IA methods (4). Rigorous, well-motivated methods (e.g., using causality) are critical, rather than correlative evidence that may not be correct or faithful. We believe that standardized, robust and widely accepted methods will increase trust in IA work, and lead to the easier and wider adoption of IA methods.
Due to the constraints of space and time, we note that it would be difficult for one work to address all these points while still making a focused contribution. Thus, we stress that our call to action is for IA research as a whole to revisit its priorities, rather than a checklist for individual papers to address.
@@
8 Related work
The increasing number of IA publications during the last few years has resulted in several survey or position papers that critically discuss existing work, identify common patterns, and provide suggestions for how to go forward. Lipton (2018) critically questions common motivations behind interpretability and the lack of definitions in the field. Following their recommendation, we provide a definition of what we consider interpretability and analysis research in §2. Belinkov and Glass (2019) summarize trends in early IA work and discuss recommendations for how to overcome the limitations of IA research. Similar to our work, they recommend that future work should think about better ways to evaluate IA research and findings. Rogers et al. (2020) survey and synthesize IA work on BERTology, a subfield of IA work that focuses on encoder-only language models. Rauker et al. (2023) survey a large number of papers that study the internals of language models (transparency), and discuss key challenges in the field. Like us, they also argue for better ways of evaluating IA methods, as well as more actionability and grounding in real-world applications. More recently, Madsen et al. (2024) discuss two prominent trends in interpretability research (post-hoc explanations and intrinsic interpretability) and argue that interpretability (“the study of explaining models in understandable terms to humans”) needs a new paradigm centered around faithfulness.
Several other works study citational patterns and trends within the broader NLP community. Mohammad (2020) uses citations to measure the impact of NLP publications indexed by the ACL Anthology. Like us, they compare how well papers from different areas within NLP are cited, and use citation statistics to draw conclusions about the impact of different subfields within NLP. Singh et al. (2023) use citations as an indicator for how widely the community is reading. They demonstrate a recency bias in citation behavior with a study of temporal citation trends, i.e., a majority of cited papers fall within a five year time period before publication of the citing work. Wahle et al. (2023) analyze the influence between NLP and other fields over the years. Also using Semantic Scholar, they rely on citations to conclude that NLP has become more insular over time. Similarly, Subramonian et al. (2024) find low levels of extra-disciplinary citation when analyzing how NLP and ML researchers discuss democracy. More specific to IA, Jacovi (2023) uses Semantic Scholar to curate a large number of papers focusing on explainability, studying citation trends in the field based on this collection.
Another set of related papers surveys the NLP community for their perceptions and opinions, a method we also use. Gururaja et al. (2023), for example, focus on paradigm shifts and study factors that shape NLP as a field. They conduct interviews with NLP researchers and experts and gather their opinions on critical trends and patterns that emerge in the field. Pramanick et al. (2023) also focus on paradigm shifts and impact, but from a diachronic perspective. They provide a novel framework to study the evolution of research topics within a field to establish what drives research in NLP across time, and they find that tasks and methods have a bigger impact on the field than metrics do.
Lastly, there are several papers in the scientometrics literature that study and compare the impact of research using the same metrics as we do: Chacon et al. (2020) apply the citation success index to compare sub-fields in physics, and Leydesdorff (2007) propose the use of Betweenness Centrality as a measure of the interdisciplinarity of journals.
@@ 9 Conclusion
Although ACL and EMNLP are the most cited *CL venues (Mohammad, 2020), our analysis excludes several other big NLP venues, including EACL, NAACL, AACL, TACL, and BlackboxNLP, a workshop which focuses on IA work. Additionally, given the growing interest in NLP, and in particular, LLMs, from the broader machine learning community, there is an increasing number of IA papers published at machine learning conferences such as ICLR, NeurIPS, and ICML, which we also do not consider in our analyses. Similarly, a vast amount of work on mechanistic interpretability has been published as articles (e.g., on LessWrong [^6] https://www.lesswrong.com/ and the AI Alignment Forum [^7] https://www.alignmentforum.org/), and blog posts (e.g., by Anthropic [^8] https://www.anthropic.com/). Therefore, there is a risk that our analysis misses potentially influential IA work published at these venues.
@@
As our analysis focuses on papers published between 2018 and 2024, our results represent a snapshot in time on the scale of research in NLP, where models and methods come and go. The time period that we look at is dominated by transformer-based language models, and a paradigm of using large, general-purpose pre-trained models for many tasks, and thus many IA papers focus on studying these. Understanding this as the context of our analysis and results is important, as they may look completely different in a time period where the most popular models and IA methods are different. This also means that our results cannot speak to the impact of today’s IA work, which will only become clear in the future.
@@
Although our use of citations is an important component of how we quantify impact in this paper, we do not consider citational context or distinguish between types of citations. However, papers can be cited for a number of reasons (Bornmann and Daniel, 2008), not all positive and not all having to do with the conventions of scholarly publishing (Bornmann and Daniel, 2008; Zhu et al., 2015; Bornmann and Marx, 2012).
Limitations of this survey
As with all surveys, our survey results might be subject to selection bias. To mitigate this risk, we took the following steps: (1) We used public mailing lists such as corpora-list to advertise our survey outside our personal networks. (2) Our social media and academic networks are diverse as we are authors from four different institutions, covering four different continents, and we are at various career stages (Masters student, PhD candidate, postdoctoral researcher, assistant professor, and full professor). (3) We targeted 100+ survey responses (and received 138). Despite our efforts to get a large number and diversity of responses, they may not be representative of the field as a whole. In particular, full professors (N=5, at various career stages), and industry practitioners who are not researchers (N=1) were somewhat underrepresented in our responses, indicating that our results focus more on research impact rather than impact on industry applications, and are mostly shaped by PhD students (41.3% of respondents), whose interests, incentives, and assessment of impact are sure to be different from respondents at other career stages.
As for survey content, some respondents brought up the following concerns about our design choices: one respondent felt our definition of IA was too broad for their taste, but our inclusion of interpretability and analysis was by design (see Section 3). Another respondent noted that we defined IA but not what we meant by “progress,” which was also by design, as we did not want to impose a normative definition of progress on our respondents but rather, get at their own intuitions, regardless of how they might define progress. Finally, one respondent complained that our questions about the usefulness of IA (to various subfields, on one’s own research, etc.) were framed in absolute rather than relative terms, and that just because IA research has some positive impact on our understanding doesn’t mean that it is the best option to pursue given limited time and resources. This paper presents views of absolute and relative impact via the survey and citation graph analyses, for a holistic view of IA research that also allows for it to have value for its own sake. Ultimately, we believe that a view of “optimal” impact compared to other options lies in the eye of the beholder, and is one (but not the only) way of interpreting our results.
@@ Acknowledgments
TVB was funded by the Centro Nacional de Inteligencia Artificial, CENIA, FB210017, Basal ANID, MM was supported by the Mila-Samsung grant, and VG was funded by the BMBF’s (German Federal Ministry of Education and Research) SLIK project under the grant 01IS22015C.
Appendix A Citation graph details
@@
Phonology, Morphology and Word Segmentation 72
@@
Unsupervised and Weakly Supervised Methods in NLP 17
@@ A.1 Sanity checks
Citation trends of IA exclusively inside ACL and EMNLP
As some of our findings depend on labels from our classifier, which might be noisy, we verify these findings using a smaller subset of our data, consisting exclusively of ACL and EMNLP papers. This subgraph of our citation graph has gold labels for the submission track. Specifically, we verify that (1) IA work is primarily cited by tracks other than IA (see Figure 4), and that (2) there is significant variation in how frequently different tracks cite IA work (see Figure 11).
In our gold labeled subgraph, there are 2,283 citations to IA papers, of which only 846 are from other IA papers (37.1%). This shows that a large fraction of citations to IA papers do indeed come from outside IA research. This verifies our first classifier-based result.
Next, when looking at the references of papers in each submission track within our gold subgraph, we find that the proportion of references to IA papers does indeed differ considerably by track. As an example, for the Large Language Models track, we find that 11.2% (N=723) of its references are to IA papers. In contrast, only 1.3% (N=1828) of the references of Sentiment Analysis track papers correspond to IA work. This confirms our second classifier-based finding.
@@
Figure 8: Betweenness centralities versus citation counts for papers in ACL and EMNLP since 2020. No correlation can be detected to the naked eye between these metrics.
@@
Appendix B Survey details
@@
B.1 Ethical considerations
Our survey involved research with human participants, thus we report the full text of the survey below, and information about recruitment in Section 3. We determined there to be a negligible risk of harms from participating in our survey, as it contains no offensive or harmful content. As shown in the full survey below, we describe our study objectives and remind respondents that filling out the survey is completely voluntary. We then explicitly ask for their consent to participate, and obtain consent from all 138 survey respondents. For respondents who may not have completed the survey, no data was collected. In lieu of financial compensation, we offered survey respondents the optional opportunity to provide their name or an alias that we would mention in the acknowledgements of any future paper we write with the survey results. To protect respondent privacy and confidentiality, when we report disaggregated results in this paper, we ensured a minimum of 10 respondents per bucket. In addition, we will not release the original survey responses in full, but only release high-level statistics, annotations from our qualitative coding, and select non-identifying examples in Section 7.
B.2 Participant demographics
We collected demographic information (occupations and research areas) from survey respondents to consider factors that might affect the representativeness of our results. Table 4 presents the breakdown of respondents per occupation.
Occupation Responses
PhD student/candidate 57 (041%)
Postdoc 15 (011%)
Junior industry researcher 17 (012%)
Master’s student 12 (008%)
Assistant professor 10 (007%)
Senior industry researcher 10 (007%)
Bachelor’s student 6 (005%)
Full professor 5 (004%)
Associate professor 2 (001%)
NLP Practitioner 1 (001%)
Other (write-in) 3 (002%)
Total 138 (100%)
Table 4: Raw numbers and percentages of survey responses, grouped by the respondent’s occupation.
When collecting information on research areas, we allowed respondents to check multiple boxes corresponding to multiple research areas. Of particular interest is the area labeled "Science of LMs," which we used as an umbrella term to include analysis and interpretability research. Table 5 shows the research areas of our respondents. The next section provides the expansions for each of the umbrella terms that we list in the table.
Research area Responses
Science of LMs 54 (39%)
Evaluation 53 (38%)
LM adaptation 47 (34%)
Data for LMs 32 (23%)
NLP applications 32 (23%)
Computational linguistics 30 (22%)
Mind, brain and LMs 30 (22%)
Neurosymbolic approaches 26 (19%)
Learning algorithms 25 (18%)
LMs for everyone 24 (17%)
LMs and the world 21 (15%)
Safety 21 (15%)
Societal implications 20 (14%)
Inference algorithms 14 (10%)
Multimodal and novel applications 14 (10%)
Compute-efficient LMs 10 (07%)
Table 5: Raw numbers and percentages of survey respondents who selected a certain research area. Respondents were allowed to select multiple areas, which is why the numbers add up to more than 138. Refer to the full survey for details on what each umbrella term represents.
B.3 Full survey
@@ 9. Do you think model analysis and interpretability research is important, and if so, why?
Appendix C Qualitative coding
@@
Appendix D Additional results
@@
Figure 10: Betweenness centrality of ACL and EMNLP papers since 2020 by track. Lines at the middle of the box represent the medians, but some tracks have their median at 0. IA papers are more central than papers from most tracks.
@@
Figure 11: Percentage of references to IA papers according to our classifiers prediction for different tracks. There are significant differences across tracks in how IA is cited. This is also true when only considering gold labels for tracks (see Section A.1).
Figure 12: Ratio of intra-track citations according to the predictions of our classifier. It measures the percentage of citations to papers of track A from papers that are also in track A. IA does not stand out in terms of the percentage of citations which are made by other papers of its own track.